If the price of a slice of pizza rises from $2.50 to $3, and quantity demanded falls

from 10,000 slices to 7,400 slices, using the formula for arc price elasticity (5 marks)

i. What is the percentage change in price?

ii. What is the percentage change in quantity?

iii. What is the level of elasticity?

iv. What is the interpretation the result


1
Expert's answer
2021-07-27T08:55:01-0400

i.) percentage change in price


"=change in price\/original" "price*100"

=0.5/2.5*100

=20%


ii.) the percentage change in quantity


= "change in quantity\/initial quantity*100"

=2600/10000*100

=26%


iii.)  the level of elasticity

="percentage change in quantity\/percentage change in price"

=26/20

=1.3


iv.)The demand is elastic since the price elasticity of demand is greater than one.